# Generalising Traffic Forecasting to Regions without Traffic Observations

### v1

@@ Results
Table Implementation Details. reports forecast errors averaged over two hours (24 time steps for highway datasets, and 8 time steps for urban traffic datasets). Our model, with either variant GenCast-H (using GeoHash embeddings) or GenCast-L (using LLM embeddings), consistently outperforms all competitors. Our model reduces forecast errors by up to 3.1% and increases $\mathrm{R}^{2}$ by up to 125.6% on the Melbourne dataset. We further conducted paired t-tests and Wilcoxon signed-rank tests between our model and the best baseline across all datasets. The results show that our model consistently outperforms the baselines with statistically significant improvements $p\ll 10^{-8}$.
@@
As shown in Table Implementation Details., GenCast-L achieves better performance on PEMS07 and PEMS08, while GenCast-H performs better on PEMS-Bay, METR-LA, and Melbourne. We attribute this discrepancy to the quality of SE-L. The data for PEMS-Bay and METR-LA were collected a long time ago (see Table 3 in Appendix C.1), whereas the regional information used for SE-L is retrieved from the latest OpenStreetMap data. Changes in the physical environment over time may result in a mismatch between the generated embeddings and the conditions at the data collection time. In addition, the Melbourne dataset covers a small, homogeneous area concentrated in Melbourne CBD, making it difficult for SE-L to capture distinctive spatial features or meaningful propagation patterns. In contrast, SE-H is generated from geo-coordinates and is updated during training, making it more adaptable and robust to different environments.
@@
[…]
Table 2: Performance on PEMS-Bay (Ring Split) and NREL.
@@
We vary the unobserved ratio (i.e., percentage of unobserved nodes in $G$) from 0.2 to 0.8 on all datasets. As before, we split each dataset either horizontally or vertically and report the performance averaged over four setups. We plot the top-3 best baselines with our models in Fig. 6. GenCast consistently performs the best, confirming its robustness against the unobserved ratio.
@@
We further adapt GenCast to the solar power NREL dataset (NREL 2018), using GenCast-w/wx (i.e., without spatial embeddings or physics constraints) due to domain complexity. As shown in Table 2, it still outperforms all baselines, demonstrating strong generalisability.
@@ Conclusion
We proposed a model named GenCast to address the challenges in traffic forecasting for unobserved regions. Unlike purely data-driven models, GenCast uses physics knowledge and external spatial-temporal data (i.e., weather) to achieve generalisability over regions without traffic observations. To support physics-informed learning, we designed two continuously differentiable spatial embeddings. We further introduced a spatial grouping module to filter out localised features that are not transferable to unobserved regions. We evaluated GenCast on real-world traffic datasets. The results show that it consistently outperforms SOTA models across different settings, achieving up to 3.1% reduction in forecast error and up to 125.6% improvement in $\mathrm{R}^{2}$.
@@ Appendix A A Related Work
Extrapolation, kriging, and forecasting are core spatial-temporal tasks that often share backbone spatial-temporal learning models (Hu et al. 2024b, c), with recent approaches combining graph and sequence models to capture spatial-temporal dependencies (Wu et al. 2019; Zheng et al. 2020; Hu et al. 2023, 2024a). Kriging and extrapolation are particularly challenging due to the lack of ground truth at unobserved target locations. To simulate such settings during training, existing models typically mask observed locations (Wu et al. 2021; Zheng et al. 2023) or interpolate between them (Xu et al. 2025), and are trained to reconstruct the masked values. While being effective for scattered missing observations, these approaches often fail in more realistic scenarios involving large, continuous unobserved regions. To address this issue, STSM (Su et al. 2024b) introduces a selective masking strategy based on location similarity, encouraging the model to learn from locations that resemble those in the unobserved region. However, because the similarity is derived from static auxiliary features (e.g., POI categories and geographic coordinates), it fails to capture dynamic traffic patterns, thereby limiting the STSM’s generalisability in complex real-world environments.
@@
CL-based models learn transferable representations by aligning similar inputs while distinguishing dissimilar ones through positive and negative sample pairs. For spatial-temporal forecasting, models often construct different graph views by perturbing nodes or edges, helping to uncover invariant patterns across space and time (Liu et al. 2022). To enable generalisation to unobserved regions, STSM (Su et al. 2024b) contrasts the forecasts learned for the full input graph with those learned for selectively masked graphs, encouraging consistency under simulated unobserved scenarios. We also follow a contrastive learning backbone. Instead of selective masking, we adopt a more flexible random subgraph masking strategy, which does not require knowledge about the similarity between the locations in the observed and the unobserved regions as mentioned above.
@@
Models in the first category leverage neural Ordinary Differential Equations (ODEs) or diffusion-inspired dynamics to simulate latent physical processes (Hwang et al. 2021; Ji et al. 2022; Wang et al. 2022; Verma, Heinonen, and Garg 2024). Such models embed physics priors implicitly through model design, capturing continuous-time evolution aligned with domain dynamics. They require fully observed spatial domains to initialise physical states, limiting their applicability to settings with unobserved regions.
Models in the second category explicitly incorporate physical constraints into the training objective. Physics-Informed Neural Networks (PINNs) (Raissi, Perdikaris, and Karniadakis 2019) extend this idea by introducing residuals of governing equations directly into the loss function, enabling models to learn solutions consistent with physical laws. PINNs have been successfully applied to domains such as air quality estimation (Hettige et al. 2024) and climate dynamics (Hwang et al. 2021), where robustness and generalisation with sparse or noisy observations are critical. For traffic forecasting, classical physical models such as the Lighthill-Whitham-Richards (Lighthill and Whitham 1955; Richards 1956) and Greenshields (Greenshields 1935) equations define macroscopic relationships among speed, density, and flow, and have been adopted as soft constraints (Shi, Mo, and Di 2021; Zhang et al. 2024).
@@
Beyond traffic observations, a variety of external (auxiliary) signals have been explored to enhance robustness and generalisation in spatial-temporal forecasting tasks. These signals can be broadly grouped into: (1) multi-sourced spatial-temporal data e.g., traffic flow data, taxi demand data, and traffic index statics (Li et al. 2024); (2) external contextual features, e.g., weather conditions, major events, and holidays (Su et al. 2024a). Most existing work focuses on using these signals to enhance sequence-based models to detect irregular events or handle short-term missing data. The use of such signals for guiding model generalisation to unobserved regions, especially when combined with advanced spatial-temporal graph networks, remains underexplored.
@@ B.1 Details of GenCast Backbone
Following Su et al. (2024b), our model backbone takes a contrastive learning architecture (Fig. 7). At each training epoch, a masked graph view $G_{o}^{m}$ is constructed by randomly masking a subgraph from the observed graph $G_{o}$, yielding two input sequences (i.e., two views): the original $\mathbf{X}_{G_{o}}^{t-T+1:t}$ and the masked $\mathbf{X}_{G_{o}^{m}}^{t-T+1:t}$. Then, we generate pseudo-observations for the masked locations, allowing the construction of both spatial (i.e., geo-coordinate) proximity-based and temporal (i.e., traffic series) similarity-based adjacency matrices.
@@
The spatial adjacency matrix $\textbf{A}_{sg}$ is defined by Eq. 13, where $\epsilon_{sg}$ is a hyperparameter (value set following Su et al. (2024b)), and $dist(c_{i},c_{j})$ denotes the Euclidean distance between locations $i$ and $j$ ($c_{i}$ and $c_{j}$ are their geo-coordinates.)
@@
GenCast uses Dynamic Time Warping (DTW) (Berndt and Clifford 1994) to compute the temporal similarity-based (non-symmetric) adjacency matrix, which links $q_{kk}$ most similar observed–observed location pairs and $q_{ku}$ most similar observed–unobserved (masked) pairs. GenCast constructs directed edges from observed to unobserved nodes only, to prevent unobserved representations from affecting observed node embeddings (by message passing). The temporal adjacency matrix for model training is denoted as $\mathbf{A}_{{dtw}}^{{train}}\in\mathbb{R}^{N_{o}\times N_{o}}$, and that for model testing is $\mathbf{A}_{{dtw}}\in\mathbb{R}^{N\times N}$. As masking is dynamic across epochs, $\mathbf{A}_{{dtw}}^{{train}}$ is updated at each training epoch.
@@
The final loss combines forecast error (RMSE) and representation alignment: $L=L_{pred}+\lambda L_{cl}$, where $\lambda$ is a weighting factor. We follow the default hyperparameter from Su et al. (2024b) without fine-tuning in GenCast.
@@ B.2 Spatial Embeddings of GenCast
Generating location descriptions is the first step in our LLM-based spatial embedding module. We construct a textual description for each location by incorporating four categories of information: (i) the location’s geo-coordinates, including latitude, longitude, and address; (ii) geometric properties describing the shape and extent of the surrounding area, particularly the spatial coverage of nearby POIs; (iii) POI information (Su et al. 2024b), such as the categories and counts of POIs within the region; and (iv) attributes of the closest road segments, including road type, maximum speed, number of lanes, and whether the road is one-way. Fig. 9 shows an example of a location description used as a prompt for generating spatial embeddings via a large language model (LLaMA 3 8B Instruct (Meta 2024)). The dimension of each spatial embedding generated by the LLM is 4,096.
@@
B.3 Physics-informed Module in GenCast
The physics-informed module introduces a constraint based on physical principles of traffic systems formulated by the Lighthill–Whitham–Richards (LWR) model (Lighthill and Whitham 1955; Richards 1956). LWR is a macroscopic first-order traffic flow model that describes the evolution of traffic density over space and time using a conservation law:
@@ B.3 Physics-informed Module in GenCast
As benchmark datasets often lack traffic density observations, we rewrite the formulation using velocity, assuming a closed system with a functional density–velocity relationship (Greenshields 1935):
@@
Here, $x_{fspd}$ denotes Free-Flow Traffic Speed, which refers to the speed at which vehicles travel under low traffic density (i.e., no delays caused by congestion, signals, or other disruptions). Following Xiong, Zhou, and Bennett (2023), we define the 85th percentile of traffic speed during non-peak hours (i.e., 9:00am–4:00pm and 10:00pm–6:00am) as the free flow speed for each location, i.e., $\mathbf{X}_{i,{fspd}}=\mathrm{Percentile}_{85}\left(\mathbf{X}_{i}^{t}\mid t\in\mathcal{T}_{{non-peak}}\right),\forall i\in G_{o}]$. We estimate $\mathbf{X}_{i,{fspd}}$ for all observed locations from the training set. The spatial and temporal derivatives hence can be rewritten as:
@@
To guide GenCast to follow the physical principles, we use the Huber loss (Huber 1992) to minimise the physics residual (i.e., to penalise violations of the physical law):
@@ C.1 Detailed Experimental Setup
We conduct experiments on four highway traffic datasets (PEMS-Bay, PEMS07, PEMS08 (CalTrans 2001) and METR-LA) and an urban traffic dataset (Melbourne). PEMS-Bay, PEMS07, and PEMS08 contain traffic speed data collected by 358, 400 and 400 sensors in California, METR-LA (Li et al. 2018) contains traffic speed data collected by 207 sensors in Los Angeles County. All highway traffic datasets are collected at 5-minute intervals, i.e., 288 time slots per day. The urban traffic dataset Melbourne is collected by 182 sensors in Melbourne at 15-minute intervals, i.e., 96 time slots per day. Additionally, we generalise our proposed model GenCast to a solar power dataset, named NREL (NREL 2018), which contains solar power data collected by 137 sensors in Alabama. Table 3 lists the dataset statistics, and Fig. 10 visualises the sensor distribution among all datasets.
[…]
@@
[…]
@@ C.3 Ablation Study
[…]
@@ C.4 Parameter and Case Study
Figure 15: Average temporal attention to weather on the PEMS07 dataset.

### v2

@@ Results
Table 1 reports forecast errors averaged over two hours (24 time steps for highway datasets, and 8 time steps for urban traffic datasets). Our model, with either variant GenCast-H (using GeoHash embeddings) or GenCast-L (using LLM embeddings), consistently outperforms all competitors. Our model reduces forecast errors by up to 3.1% and increases $\mathrm{R}^{2}$ by up to 125.6% on the Melbourne dataset. We further conducted paired t-tests and Wilcoxon signed-rank tests between our model and the best baseline across all datasets. The results show that our model consistently outperforms the baselines with statistically significant improvements $p\ll 10^{-8}$.
@@
As shown in Table 1, GenCast-L achieves better performance on PEMS07 and PEMS08, while GenCast-H performs better on PEMS-Bay, METR-LA, and Melbourne. We attribute this discrepancy to the quality of SE-L. The data for PEMS-Bay and METR-LA were collected a long time ago (see Table 3 in Appendix C.1), whereas the regional information used for SE-L is retrieved from the latest OpenStreetMap data. Changes in the physical environment over time may result in a mismatch between the generated embeddings and the conditions at the data collection time. In addition, the Melbourne dataset covers a small, homogeneous area concentrated in Melbourne CBD, making it difficult for SE-L to capture distinctive spatial features or meaningful propagation patterns. In contrast, SE-H is generated from geo-coordinates and is updated during training, making it more adaptable and robust to different environments.
@@
Table 2: Results on PEMS-Bay (Ring Split) and NREL.
@@
We vary the ratio of unobserved nodes in $G$ from 0.2 to 0.8 on all datasets. Each dataset is split horizontally or vertically, and results are averaged over four setups. Fig. 5 compares the top three baselines with our model. GenCast consistently performs the best, confirming its robustness against the unobserved ratio.
@@
We further adapt GenCast to the solar power NREL dataset (nrel), using GenCast-w/wx (i.e., without spatial embeddings or physics constraints) due to domain complexity. As shown in Table 2, it still outperforms all baselines, demonstrating strong generalisability.
@@ Conclusion
We proposed a model named GenCast to address the challenges in traffic forecasting for unobserved regions. Unlike purely data-driven approaches, GenCast uses physics knowledge and external spatial-temporal data (e.g., weather) to enhance generalisation to unobserved regions. It employs two continuously differentiable spatial embeddings to support physics-informed learning. A spatial grouping module filters localised, non-transferable features. We evaluated GenCast on real-world traffic datasets. The results show that it consistently outperforms SOTA models across different settings, achieving up to 3.1% reduction in forecast error and up to 125.6% improvement in $\mathrm{R}^{2}$. Although GenCast achieves strong generalisation across regions, the relatively low $\mathrm{R}^{2}$ indicates room for improvement. Future work will focus on balancing cross-region generalisability and local fidelity through adaptive or hybrid modelling strategies.
Ethical Statement
All datasets used in this study are publicly available and do not contain any personally identifiable information. There are no ethical concerns associated with this work.
Acknowledgments
This work is in part supported by the Australian Research Council (ARC) via Discovery Projects DP230101534 and DP240101006. Jianzhong Qi is supported by ARC Future Fellowship FT240100170. Feng Liu is supported by the ARC with grant numbers DE240101089, LP240100101, DP230101540 and the NSF&CSIRO Responsible AI program with grant number 2303037.
@@ Appendix A A Related Work
Extrapolation, kriging, and forecasting are core spatial-temporal tasks that often share backbone spatial-temporal learning models (ustd; stgp), with recent approaches combining graph and sequence models to capture spatial-temporal dependencies (graphwavenet; gman; STGNP; dualSTN). Kriging and extrapolation are particularly challenging due to the lack of ground truth at unobserved target locations. To simulate such settings during training, existing models typically mask observed locations (ignnk; increase) or interpolate between them (kits), and are trained to reconstruct the masked values. While being effective for scattered missing observations, these approaches often fail in more realistic scenarios involving large, continuous unobserved regions. To address this issue, STSM (stsm) introduces a selective masking strategy based on location similarity, encouraging the model to learn from locations that resemble those in the unobserved region. However, because the similarity is derived from static auxiliary features (e.g., POI categories and geographic coordinates), it fails to capture dynamic traffic patterns, thereby limiting the STSM’s generalisability in complex real-world environments.
@@
CL-based models learn transferable representations by aligning similar inputs while distinguishing dissimilar ones through positive and negative sample pairs. For spatial-temporal forecasting, models often construct different graph views by perturbing nodes or edges, helping to uncover invariant patterns across space and time (whdocl). To enable generalisation to unobserved regions, STSM (stsm) contrasts the forecasts learned for the full input graph with those learned for selectively masked graphs, encouraging consistency under simulated unobserved scenarios. We also follow a contrastive learning backbone. Instead of selective masking, we adopt a more flexible random subgraph masking strategy, which does not require knowledge about the similarity between the locations in the observed and the unobserved regions as mentioned above.
@@
Models in the first category leverage neural Ordinary Differential Equations (ODEs) or diffusion-inspired dynamics to simulate latent physical processes (hwang2021climate; STDEN; st-pef; ClimODE). Such models embed physics priors implicitly through model design, capturing continuous-time evolution aligned with domain dynamics. They require fully observed spatial domains to initialise physical states, limiting their applicability to settings with unobserved regions.
Models in the second category explicitly incorporate physical constraints into the training objective. Physics-Informed Neural Networks (PINNs) (raissi2019physics) extend this idea by introducing residuals of governing equations directly into the loss function, enabling models to learn solutions consistent with physical laws. PINNs have been successfully applied to domains such as air quality estimation (hettigeairphynet) and climate dynamics (hwang2021climate), where robustness and generalisation with sparse or noisy observations are critical. For traffic forecasting, classical physical models such as the Lighthill-Whitham-Richards (lwr1; lwr2) and Greenshields (greenshields) equations define macroscopic relationships among speed, density, and flow, and have been adopted as soft constraints (shi2021physics; zhang2024physics).
@@
Beyond traffic observations, a variety of external (auxiliary) signals have been explored to enhance robustness and generalisation in spatial-temporal forecasting tasks. These signals can be broadly grouped into: (1) multi-sourced spatial-temporal data e.g., traffic flow data, taxi demand data, and traffic index statics (li2024opencity); (2) external contextual features, e.g., weather conditions, major events, and holidays (dualcast). Most existing work focuses on using these signals to enhance sequence-based models to detect irregular events or handle short-term missing data. The use of such signals for guiding model generalisation to unobserved regions, especially when combined with advanced spatial-temporal graph networks, remains underexplored.
@@ B.1 Details of GenCast Backbone
Following stsm, our model backbone takes a contrastive learning architecture (Fig. 6). At each training epoch, a masked graph view $G_{o}^{m}$ is constructed by randomly masking a subgraph from the observed graph $G_{o}$, yielding two input sequences (i.e., two views): the original $\mathbf{X}_{G_{o}}^{t-T+1:t}$ and the masked $\mathbf{X}_{G_{o}^{m}}^{t-T+1:t}$. Then, we generate pseudo-observations for the masked locations, allowing the construction of both spatial (i.e., geo-coordinate) proximity-based and temporal (i.e., traffic series) similarity-based adjacency matrices.
@@
The spatial adjacency matrix $\textbf{A}_{sg}$ is defined by Eq. 13, where $\epsilon_{sg}$ is a hyperparameter (value set following stsm), and $dist(c_{i},c_{j})$ denotes the Euclidean distance between locations $i$ and $j$ ($c_{i}$ and $c_{j}$ are their geo-coordinates.)
@@
GenCast uses Dynamic Time Warping (DTW) (dtw) to compute the temporal similarity-based (non-symmetric) adjacency matrix, which links $q_{kk}$ most similar observed–observed location pairs and $q_{ku}$ most similar observed–unobserved (masked) pairs. GenCast constructs directed edges from observed to unobserved nodes only, to prevent unobserved representations from affecting observed node embeddings (by message passing). The temporal adjacency matrix for model training is denoted as $\mathbf{A}_{{dtw}}^{{train}}\in\mathbb{R}^{N_{o}\times N_{o}}$, and that for model testing is $\mathbf{A}_{{dtw}}\in\mathbb{R}^{N\times N}$. As masking is dynamic across epochs, $\mathbf{A}_{{dtw}}^{{train}}$ is updated at each training epoch.
@@
The final loss combines forecast error (RMSE) and representation alignment: $L=L_{pred}+\lambda L_{cl}$, where $\lambda$ is a weighting factor. We follow the default hyperparameter from stsm without fine-tuning in GenCast.
@@ B.2 Spatial Embeddings of GenCast
Generating location descriptions is the first step in our LLM-based spatial embedding module (sun2025flexireg). We construct a textual description for each location by incorporating four categories of information: (i) the location’s geo-coordinates, including latitude, longitude, and address; (ii) geometric properties describing the shape and extent of the surrounding area, particularly the spatial coverage of nearby POIs; (iii) POI information (stsm; hafusion), such as the categories and counts of POIs within the region; and (iv) attributes of the closest road segments, including road type, maximum speed, number of lanes, and whether the road is one-way. Fig. 8 shows an example of a location description used as a prompt for generating spatial embeddings via a large language model (LLaMA 3 8B Instruct (llama2024)). The dimension of each spatial embedding generated by the LLM is 4,096.
@@
B.3 Spatial Grouping Module in GenCast
Fig. 9 illustrates the spatial grouping module, which softly clusters locations into latent groups, allowing GenCast to capture shared group-level patterns while suppressing ad hoc variations at individual locations through an entropy regularisation term.
B.4 Physics-informed Module in GenCast
The physics-informed module introduces a constraint based on physical principles of traffic systems formulated by the Lighthill–Whitham–Richards (LWR) model (lwr1; lwr2). LWR is a macroscopic first-order traffic flow model that describes the evolution of traffic density over space and time using a conservation law:
@@ B.4 Physics-informed Module in GenCast
As benchmark datasets often lack traffic density observations, we rewrite the formulation using velocity, assuming a closed system with a functional density–velocity relationship (greenshields):
@@
Here, $x_{fspd}$ denotes Free-Flow Traffic Speed, which refers to the speed at which vehicles travel under low traffic density (i.e., no delays caused by congestion, signals, or other disruptions). Following fspd, we define the 85th percentile of traffic speed during non-peak hours (i.e., 9:00am–4:00pm and 10:00pm–6:00am) as the free flow speed for each location, i.e., $\mathbf{X}_{i,{fspd}}=\mathrm{Percentile}_{85}\left(\mathbf{X}_{i}^{t}\mid t\in\mathcal{T}_{{non-peak}}\right),\forall i\in G_{o}]$. We estimate $\mathbf{X}_{i,{fspd}}$ for all observed locations from the training set. The spatial and temporal derivatives hence can be rewritten as:
@@
To guide GenCast to follow the physical principles, we use the Huber loss (huber1992robust) to minimise the physics residual (i.e., to penalise violations of the physical law):
@@ C.1 Detailed Experimental Setup
We conduct experiments on four highway traffic datasets (PEMS-Bay, PEMS07, PEMS08 (pems) and METR-LA) and an urban traffic dataset (Melbourne). PEMS-Bay, PEMS07, and PEMS08 contain traffic speed data collected by 358, 400 and 400 sensors in California, METR-LA (dcrnn) contains traffic speed data collected by 207 sensors in Los Angeles County. All highway traffic datasets are collected at 5-minute intervals, i.e., 288 time slots per day. The urban traffic dataset Melbourne is collected by 182 sensors in Melbourne at 15-minute intervals, i.e., 96 time slots per day. Additionally, we generalise our proposed model GenCast to a solar power dataset, named NREL (nrel), which contains solar power data collected by 137 sensors in Alabama. Table 3 lists the dataset statistics, and Fig. 10 visualises the sensor distribution among all datasets.
  Dataset Data Type Time period Interval #Sensors
@@
  Model Time PEMS-Bay PEMS07 PEMS08 Melbourne METR
@@ C.3 Ablation Study
  Dataset Metric GenCast-H GenCast-H (L2) GenCast-L GenCast-L (L2)
@@ C.4 Parameter and Case Study
Figure 15: Average temporal attention to weather on the PEMS07.